Circular No. 7869

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 AR_{2} (LINEAR) An apparently asteroidal object reported by LINEAR in early Jan-

uary (discovery position below) and considered of only moderate interest was also recognized at the Minor Planet Center to have been accidentally reobserved by LINEAR on three isolated nights in February and March, after which the orbit calculation suggested the object might be a comet. It was therefore listed on the NEO Confirmation Page. Observations with the 1.2-m reflector at Mt. Hopkins Observatory by P. Berlind and T. B. Spahr on Apr. 6.1 UT in poor conditions (high thin cirrus) showed that the images were softer than those of stars of similar brightness, and observations on Apr. 7.2 showed the slightly diffuse images of a definite comet.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 6.14531  2^{h}32^{m}53.^{s}60  +46^{o}57′58′′.1  18.9

Complete astrometry is tabulated on MPEC 2002-G30, together with the following orbital elements and an ephemeris:

Epoch = 2002 Jan. 6.0 TT

T = 2002 Jan. 16.5339 TT  ω = 73.^{o}6733
e = 0.615703  Ω = 7.8464 }2000.0
q = 2.064463 AU  i = 21.1092

a = 5.372055 AU n^{o} = 0.0791577 P = 12.451 years

COMET C/2002 F1 (UTSUNOMIYA) Ephemeris continuation to IAUC 7855 (MPEC 2002-G11 elements):

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 6 23^{h}23.^{m}36 +23^{o}26.^{′}3 1.255 0.610 28.^{o}7 51.^{o}8 6.8
8 23 40.05 +25 33.2 1.231 0.578 27.7 53.5 6.6
10 23 58.35 +27 32.8 1.210 0.548 26.6 55.0 6.3
12 0 18.28 +29 20.5 1.195 0.520 25.5 56.1 6.0
14 0 39.77 +30 51.4 1.184 0.495 24.4 56.8 5.8
16 1 02.57 +32 00.3 1.179 0.473 23.3 57.1 5.6
18 1 26.27 +32 42.7 1.178 0.456 22.4 56.8 5.4
20 1 50.33 +32 55.8 1.183 0.445 21.6 56.1 5.3
22 2 14.08 +32 38.7 1.192 0.439 20.9 54.8 5.3
24 2 36.92 +31 53.2 1.205 0.439 20.5 53.2 5.3
26 2 58.34 +30 42.9 1.223 0.446 20.2 51.3 5.4
2002 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden